OpenStudy (wintersuntime):

How can I find the x-intercept in the function : -x^2+1

OpenStudy (r.hamza17):

=-x^2+1 -x^2=-1 x^2=1 \[\sqrt{x^{2}}=\sqrt{1}\] x=1

OpenStudy (wintersuntime):

thank you. do you know how to find the ned behavior for it?

OpenStudy (wintersuntime):

end behavior

OpenStudy (r.hamza17):

End behavior? I didnt really get that term. Are you using the right term?

OpenStudy (r.hamza17):

Oh sorry i got ya what you were asking.

OpenStudy (wintersuntime):

ok

OpenStudy (r.hamza17):

To determine the end behavior of an equation the degree and the leading coefficient helps us. A graph is plotted and by checking the leading coefficient in our case which is negative in -x^2, so the end behavior of the function would be \[F(X)\rightarrow-\infty, as X \rightarrow-\infty\]\[F(X)\rightarrow-\infty, as X \rightarrow+ \infty \]

OpenStudy (wintersuntime):

thank u